# To curve, or not to curve: Is curvature-assisted quintessence observationally viable?

### v1

@@ 1 Introduction
In the rest of the paper, we investigate in detail the results of Ref. Andriot et al. (2023), where the authors argue that the aforementioned obstacles to finding a dynamical mechanism for cosmic acceleration and/or cosmic inflation in asymptotic regions of moduli space can be overcome if the Universe has a negative spatial curvature—the reason is the appearance of a fixed point in the phase space of the model that does not exist in the zero-curvature scenario. The arguments of Ref. Andriot et al. (2023) are based on a phase space analysis of a scalar field model with an exponential potential and in the absence of matter and radiation. Here, we analyze the model through a more realistic setup and by comparing its cosmological predictions with observational data. We show, after a qualitative discussion of the phase space of the model and through both a numerical grid analysis and a full Bayesian statistical exploration of the parameter space of the model, that allowing the curvature parameter to take nonzero values does not help and the resulting “curvature-assisted” quintessence does not provide observationally viable cosmological solutions if the potential is steep. We place an upper bound on the slope of the scalar potential that is way too low for a single-field quintessence model to satisfy the theoretical de Sitter constraint. We also show, based on qualitative arguments, that the same conclusion holds for cosmic inflation.
@@ 2 String theory and scalar fields
There is a large body of string theory literature devoted to charting the moduli space and the associated potential $V(\vec{\phi})$. The moduli space is often divided into the “bulk” region, where all moduli have a vacuum expectation value (vev) of $\mathcal{O}(1)$, and the “asymptotic” regions, where one or more moduli have a very large vev. Very little is known about the scalar potential in the bulk of moduli space, where there is no control parameter and the theory becomes subject to quantum corrections and non-perturbative effects. In contrast, in asymptotic limits of moduli space, oftentimes we can do a perturbative asymptotic expansion in some of the moduli $\phi^{a}$, and consequently $V(\vec{\phi})$ can be reliably computed. Because of this, there has been a large effort to find de Sitter minima in the asymptotic region of moduli space. However, despite decades of effort, realizing a de Sitter minimum in this asymptotic region remains an open question. In fact, every time we compute an asymptotic potential in string theory, it is a sum of exponentials,
$V=V_{0}e^{-\vec{\lambda}\cdot{\vec{\phi}}}\,,$ (1)
@@
The gradient bound in (2) implies that $V(\vec{\phi})$ is way too steep to accommodate eternally accelerating cosmological solutions (see also Refs. Agrawal et al. (2018); Akrami et al. (2019); Raveri et al. (2019); Akrami et al. (2021) for observational bounds on the slope of the potential); therefore, not even a quintessence scenario can be realized in asymptotic corners of moduli space, at least in a naive way. The situation is summarized in fig. 1—there is a bulk moduli space that we do not understand very well, where in principle, quintessence or $\Lambda$ could lurk, and the asymptotic region, which we understand well, but does not accommodate accelerated expansion.
@@
In short, from a string theory point of view, it looks like we should leave the easy asymptotic region and start working our way towards the bulk (this is indeed what most proposed de Sitter constructions do; see Refs. Kachru et al. (2003); Danielsson and Dibitetto (2013); Blåbäck et al. (2013); Kachru et al. (2021); Bena et al. (2023); Cicoli et al. (2024) and references therein). However, the recent work Andriot et al. (2023) revived the interest in asymptotic regions in string theory, showing that one can get acceleration even when $\lambda>\sqrt{2}$, simply by allowing the Universe to have a negative spatial curvature. This fact was already known in the cosmology literature van den Hoogen et al. (1999); Gosenca and Coles (2016), but Ref. Andriot et al. (2023) emphasized that this would revive the hope that one could obtain accelerated expansion in a fully controlled string theoretic EFT. In the next section, we review this construction in some detail, pointing out a number of features that significantly challenge its cosmological viability.
@@ 3 Curvature to the rescue?
We now describe the dynamical system associated with a universe filled with a single-field quintessence as dark energy, following Refs. van den Hoogen et al. (1999); Gosenca and Coles (2016); Andriot et al. (2023). We allow the spatial curvature to take a nonzero value, and we also include non-relativistic matter in the analysis with the intention of comparing the system to observational constraints. Although in a fully realistic scenario radiation should also be included, we omit its contribution to the energy budget of the Universe as it becomes negligible long before the accelerated expansion phase begins. We use this approximate dynamical system to draw qualitative conclusions about the viability of quintessence dark energy, before constraining the model in the next section through a full statistical analysis of its parameter space with no approximation.
@@ 3.1 Dynamical system
$\Omega_{\rm{m}}\equiv\frac{\rho_{m}}{3M_{\rm{Pl}}^{2}H^{2}}\,.$ (13)
@@
Based on the results of Refs. van den Hoogen et al. (1999); Gosenca and Coles (2016), it was emphasized in Ref. Andriot et al. (2023) that solving these equations could lead to viable cosmological histories with accelerated expansion for $\lambda>\sqrt{2}$ when $k=-1$ (or $\Omega_{k}>0$). This sprung much excitement in the string theory community as these are the kinds of exponential potentials that are easily encountered in string theoretic EFTs. The hope was that this could allow us to embed a realistic cosmology in the asymptotic limit of moduli space, where string theory is under perturbative control. Indeed, in Ref. Andriot et al. (2023) it was argued that one could obtain an infinite number of $e$-folds of accelerated expansion along some trajectories in phase space, with $\lambda>\sqrt{2}$. Although these statements are true, we now point out a couple of reasons why this scenario might not be as appealing as it appears at first sight.
@@ 3.2 Quality of accelerating phases
Fig. 2 shows the phase space of the dynamical system corresponding to eqs. 18, 19 and 20 for a hyperbolic (or open) universe (i.e., for $k=-1$). The paraboloid in the figure corresponds to the Friedmann equation (14) with $\Omega_{\rm{m}}=0$, above which all possible cosmic trajectories with nonzero $\Omega_{\rm{m}}$ are located. The red and the green dots show the fixed points of the system that can potentially provide trajectories with desired properties, and correspond, respectively, to unstable and attractor solutions. The figure also shows a trajectory for $\lambda=1.6$ (black curve) as an indicative example—the attractor point associated with the trajectory behaves like a node. For a more complete discussion of the fixed points and their stability see Refs. Gosenca and Coles (2016); Andriot et al. (2023).
Figure 2: Phase space of the dynamical system described by eqs. 18, 19 and 20 for a hyperbolic (or open) universe (i.e., for $k=-1$). Trajectories that pass through the green areas experience accelerating expansion with $w_{\rm{eff}}\in[-1,-0.7]$, while the red regions correspond to $w_{\rm{eff}}=0$; see eq. 16. The red and green dots show the relevant fixed points of the system and correspond, respectively, to unstable and attractor solutions. For the example of $\lambda=1.6$ the attractor point behaves like a node, as shown by the indicative trajectory (black curve), where the black dot shows the initial conditions.
@@
The key point in the curvature-assisted acceleration of Ref. Andriot et al. (2023) is that the number of $e$-folds can also diverge when $1+3w_{\rm{eff}}$ approaches zero. As noted already in that reference, this is the boundary of the accelerating region—the trajectories found in Ref. Andriot et al. (2023) with curvature asymptote to fixed points living in this boundary. Although the points themselves do not support accelerated expansion, a trajectory approaching it can support an arbitrarily large number of $e$-folds. For instance, for a fixed point at $z=z_{*}$, assuming $1+3w_{\rm{eff}}\sim C(z-z_{*})$ in the vicinity of the fixed point, where $C$ is some constant, leads to $N\sim C^{-1}\log(|z-z_{*}|)$.
@@
The curvature-assisted acceleration with steep potentials is also in serious trouble when describing an inflationary phase in the early Universe. For single-field, slow-roll cosmic inflation, there is an approximate relation, Baumann (2018)
@@
Furthermore, the fixed points with $\lambda>\sqrt{2}$ that lead to an arbitrary large number of $e$-folds of accelerated expansion with curvature in Ref. Andriot et al. (2023) have $z_{*}=-\Omega_{k}$ of $\mathcal{O}(1)$ unless $\lambda$ is very close to $\sqrt{2}$. That means that for large values of $\lambda$, curvature is an $\mathcal{O}(1)$ part of the energy budget of the Universe during the eternally accelerated trajectories. This contrasts with observations, either in inflationary or current era, which pose stringent upper bounds on $\Omega_{k}$.
@@ 4.1 Numerical grid analysis of phase space
Here, we solve the system of differential equations (18)-(20) for different initial conditions, i.e., for different initial values of the three phase space variables $\{x,y,z\}$, and then study the behavior of the cosmological background solution in each case. We do this through a fine grid scan of the entire phase space contained within the paraboloid of fig. 2, defined as $x^{2}+y^{2}-z\leq 1$; see eq. 14. We compute all the trajectories for the different sets of initial conditions and keep only the ones that provide an epoch of $w_{\rm{eff}}\leq-0.7$, i.e., solutions which go through the green regions of fig. 2—we show one example of such solutions for $\lambda=1.6$ in the figure (black curve). Fig. 3 presents the time evolution of $w_{\rm{eff}}$ for all these accelerating solutions. It is clear from the figure that none of the solutions exhibit an epoch of matter domination in the past, i.e., before the accelerating phase of cosmic history.
@@
Having said that, we do find solutions that provide extended matter domination periods, as we have demonstrated in the upper panel of fig. 4 for $\lambda=1.6$ as an example—see how $\Omega_{\rm{m}}$ behaves with time (or number of $e$-folds). However, the problem with these solutions is that none of them provide an effective equation of state parameter $w_{\rm{eff}}$ less than $\sim-0.5$ at the present time (i.e., when $\Omega_{\rm{m}}\sim 0.3$). This means that we do not obtain sufficient acceleration in these cases where matter domination exists. In the lower panel of fig. 4, we present three examples of $w_{\rm{eff}}$ evolution for the curvature-assisted quintessence model, i.e., for $\lambda=0.5,1.4,1.6$, and compare them with that of a flat $\Lambda$CDM model (i.e., with $\Omega_{k}$ set to zero) that provides a good fit to the data. The figure shows that by increasing $\lambda$ it becomes more and more difficult for the quintessence model to provide a cosmic history that is compatible with cosmological observations, as the current value of $w_{\rm{eff}}$ does not reach the observationally preferred value of $\sim-0.7$, as indicated by our $\Lambda$CDM example.
@@ 4.2 Bayesian analysis and parameter estimation
We perform a Markov chain Monte Carlo (MCMC) scan of the parameter space of the quintessence model for two cases. [^3] The Mathematica codes used in our analysis will be made publicly available in a GitHub repository upon publication of the paper. We first consider a flat universe, i.e., we fix the curvature density parameter $\Omega_{k}$ to zero. We then allow $\Omega_{k}$ to vary as a free parameter. For each case, we perform two separate analyses, one by setting the initial value of the scalar field’s time derivative $\dot{\phi}_{0}$ to zero and one by varying it as a free parameter. In both cases, we fix $\phi_{0}$ (the initial value of the scalar field) to zero and vary the parameter $V_{0}$ in the potential $V(\phi)=V_{0}e^{-\lambda\phi}$, as one of the two parameters is redundant and can be replaced by the other. We impose flat (or uniform) priors on all free parameters—table 1 shows the ranges of parameters we employ in our scans. Comparing the parameter constraints we obtain for different cases will tell us whether assisting the accelerated expansion with a nonzero spatial curvature will increase the upper bound on the parameter $\lambda$ so that large values suggested by the statement (2) become observationally viable.
@@
$\Omega_{\rm{m}}$ $[0.001,1]$
$\omega_{\mathrm{b}}$ $[0.01,0.04]$
$\Omega_{k}$ $[-0.2,0.2]$
@@
Table 1: Ranges of model parameters employed in our scans. Note that $\dot{\phi}_{0}$ is fixed to zero in some of the analyses, so the presented range for $\dot{\phi}_{0}$ is only for cases where it is varied as a free parameter.
Figure 5: The $68.3\%$ and $95.5\%$ confidence regions and one-dimensional marginalized probability density functions for the quintessence model of dark energy with an exponential potential when the spatial curvature of the Universe is set to zero (i.e., $\Omega_{k}=0$). We show the results for two cases: 1) the initial value of the scalar field’s time derivative $\dot{\phi}_{0}$ is set to zero, 2) $\dot{\phi}_{0}$ is varied as a free parameter.
Figure 6: As in fig. 5, but for cases where the spatial curvature of the Universe (i.e., $\Omega_{k}$) is varied as a free parameter.
 Parameter   $\Omega_{k}=0$ (flat universe), $\dot{\phi}_{0}=0$   $\Omega_{k}=0$ (flat universe), free $\dot{\phi}_{0}$   Free $\Omega_{k}$, $\dot{\phi}_{0}=0$   Free $\Omega_{k}$, free $\dot{\phi}_{0}$
$\Omega_{\rm{m}}$ $0.3050^{+0.0784}_{-0.0591}$ $0.3269^{+0.0730}_{-0.0777}$ $0.3134^{+0.0849}_{-0.0441}$ $0.3699^{+0.0820}_{-0.0900}$
$\omega_{\mathrm{b}}$ $0.0226^{+0.0002}_{-0.0001}$ $0.0226^{+0.0002}_{-0.0001}$ $0.0225^{+0.0001}_{-0.0002}$ $0.0224^{+0.0002}_{-0.0001}$
$\Omega_{k}$ N/A N/A $0.0047^{+0.0018}_{-0.0023}$ $0.0056^{+0.002}_{-0.003}$
@@
We present the results of our MCMC analysis in figs. 5 and 6, where two-dimensional $68.3\%$ and $95.5\%$ confidence regions and one-dimensional marginalized probability density functions (PDFs) for different parameters and different cases are provided. We also present in table 2 $1\sigma$ (i.e., $68.3\%$ confidence level) constraints on the parameters. Since the parameter of interest here is $\lambda$, let us compare our constraints on that parameter as given in the figures and the table for the different cases we have studied. Fig. 5 and table 2 show that cosmological observations place an upper bound of $\sim 0.6$ on $\lambda$ when $\dot{\phi}_{0}$ is fixed to zero, and an upper bound of $\sim 0.5$ when $\dot{\phi}_{0}$ is varied, at the $63.3\%$ confidence level, which are consistent with the results of Refs. Agrawal et al. (2018); Akrami et al. (2019); Raveri et al. (2019). When $\Omega_{k}$ is allowed to vary, we obtain the upper bound of $\sim 0.7$ on $\lambda$ (as given in fig. 6 and table 2), independently of whether $\dot{\phi}_{0}$ is varied or fixed to zero, even though the contours and the one-dimensional PDFs in the two cases are slightly different. It is, therefore, true that allowing $\Omega_{k}$ to vary pushes the upper bound on $\lambda$ to larger values, but the improvement is tiny and the favored values of $\lambda$ are far below the $\lambda=\sqrt{2}$ limit of the statement (2). This means that the new fixed point in the phase space of the model which arises as a result of $\Omega_{k}$ taking positive values, i.e., the Universe having a negative curvature, does not help with making $\lambda>\sqrt{2}$ consistent with observations—this is what we expected from the qualitative analyses of previous sections. It is, however, interesting to note that when $\Omega_{k}$ is allowed to vary, cosmological data seem to prefer nonzero curvature and therefore indicate a slight preference for an open universe, although the preference is only slightly more than $2\sigma$. We do not perform a model comparison here, but it is likely that the slightly larger upper bound on $\lambda$ for the curvature-assisted model compared to the flat one is due to the increase in the number of free parameters (i.e., the number of degrees of freedom) in the former case. As stated above, the results of the full MCMC analysis perfectly agree with the findings of the numerical grid analysis of section 4.1.

### v2

@@ 1 Introduction
In the rest of the paper, we investigate in detail the results of Refs. Marconnet and Tsimpis (2023); Andriot et al. (2023), where the authors argue that the aforementioned obstacles to finding a dynamical mechanism for cosmic acceleration and/or cosmic inflation in asymptotic regions of moduli space can be overcome if the Universe has a negative spatial curvature—the reason is the appearance of a fixed point in the phase space of the model that does not exist in the zero-curvature scenario. The arguments of Refs. Marconnet and Tsimpis (2023); Andriot et al. (2023) are based on a phase space analysis of a scalar field model with an exponential potential and in the absence of matter and radiation. Here, we analyze the model through a more realistic setup and by comparing its cosmological predictions with observational data. We show, after a qualitative discussion of the phase space of the model and through both a numerical grid analysis and a full Bayesian statistical exploration of the parameter space of the model, that allowing the curvature parameter to take nonzero values does not help and the resulting “curvature-assisted” quintessence does not provide observationally viable cosmological solutions if the potential is steep. This is true even if we allow for the initial velocity of the scalar field to vary, in order to explore the full phase space. We place an upper bound on the slope of the scalar potential that is way too low for a single-field quintessence model to satisfy the theoretical de Sitter constraint. We also show, based on qualitative arguments, that the same conclusion holds for cosmic inflation.
@@ 2 String theory and scalar fields
There is a large body of string theory literature devoted to charting the moduli space and the associated potential $V(\vec{\phi})$. The moduli space is often divided into the “bulk” region, where all moduli have a vacuum expectation value (vev) of $\mathcal{O}(1)$, and the “asymptotic” regions, where one or more moduli have a very large vev. Very little is known about the scalar potential in the bulk of moduli space, where there is no control parameter and the theory becomes subject to quantum corrections and non-perturbative effects. In contrast, in asymptotic limits of moduli space, oftentimes we can do a perturbative asymptotic expansion in some of the moduli $\phi^{a}$, and consequently $V(\vec{\phi})$ can be reliably computed. Because of this, there has been a large effort to find de Sitter minima in the asymptotic regions of moduli space. However, despite decades of effort, realizing a de Sitter minimum in these asymptotic regions remains an open question. In fact, every time we compute an asymptotic potential in string theory, it is a sum of exponentials,
$V(\vec{\phi})=V_{0}e^{-\vec{\lambda}\cdot{\vec{\phi}}}\,,$ (1)
@@
The gradient bound in (2) implies that $V(\vec{\phi})$ is way too steep for single-field models to accommodate eternally accelerating cosmological solutions (see also Refs. Agrawal et al. (2018); Akrami et al. (2019); Raveri et al. (2019); Akrami et al. (2021) for observational bounds on the slope of the potential); therefore, not even a quintessence scenario can be realized in asymptotic corners of moduli space, at least in a naive way. The situation is summarized in fig. 1—there is a bulk moduli space that we do not understand very well, where in principle, quintessence or $\Lambda$ could lurk, and the asymptotic region, which we understand well, but does not accommodate accelerated expansion.
@@
In short, from a string theory point of view, it looks like we should leave the easy asymptotic region and start working our way towards the bulk (this is indeed what most proposed de Sitter constructions do; see Refs. Kachru et al. (2003); Danielsson and Dibitetto (2013); Blåbäck et al. (2013); Kachru et al. (2021); Bena et al. (2023); Cicoli et al. (2024) and references therein). However, the recent work Marconnet and Tsimpis (2023); Andriot et al. (2023) revived the interest in asymptotic regions in string theory, showing that one can get acceleration even when $\lambda>\sqrt{2}$, simply by allowing the Universe to have a negative spatial curvature. This fact was already known in the cosmology literature van den Hoogen et al. (1999); Gosenca and Coles (2016), but Refs. Marconnet and Tsimpis (2023); Andriot et al. (2023) emphasized that this would revive the hope that one could obtain accelerated expansion in a fully controlled string theoretic EFT. In the next section, we review this construction in some detail, pointing out a number of features that significantly challenge its cosmological viability.
@@ 3 Curvature to the rescue?
We now describe the dynamical system associated with a universe filled with a single-field quintessence as dark energy, following Refs. van den Hoogen et al. (1999); Gosenca and Coles (2016); Marconnet and Tsimpis (2023); Andriot et al. (2023). We allow the spatial curvature to take a nonzero value, and we also include non-relativistic matter in the analysis with the intention of comparing the system to observational constraints. Although in a fully realistic scenario radiation should also be included, we omit its contribution to the energy budget of the Universe as it becomes negligible long before the accelerated expansion phase begins. We use this approximate dynamical system to draw qualitative conclusions about the viability of quintessence dark energy, before constraining the model in the next section through a full statistical analysis of its parameter space with no approximation.
@@ 3.1 Dynamical system
$\Omega_{\rm{m}}\equiv\frac{\rho_{\mathrm{m}}}{3M_{\rm{Pl}}^{2}H^{2}}\,.$ (13)
@@
Based on the results of Refs. van den Hoogen et al. (1999); Gosenca and Coles (2016), it was emphasized in Refs. Marconnet and Tsimpis (2023); Andriot et al. (2023) that solving these equations could lead to viable cosmological histories with accelerated expansion for $\lambda>\sqrt{2}$ when $k=-1$ (or $\Omega_{k}>0$). This sprung much excitement in the string theory community as these are the kinds of exponential potentials that are easily encountered in string theoretic EFTs. The hope was that this could allow us to embed a realistic cosmology in the asymptotic limit of moduli space, where string theory is under perturbative control. Indeed, in Refs. Marconnet and Tsimpis (2023); Andriot et al. (2023) it was argued that one could obtain an infinite number of $e$-folds of accelerated expansion along some trajectories in phase space, with $\lambda>\sqrt{2}$. Although these statements are true, we now point out a couple of reasons why this scenario might not be as appealing as it appears at first sight.
@@ 3.2 Quality of accelerating phases
Fig. 2 shows the phase space of the dynamical system corresponding to eqs. 18, 19 and 20 for a hyperbolic (or open) universe (i.e., for $k=-1$). The paraboloid in the figure corresponds to the Friedmann equation (14) with $\Omega_{\rm{m}}=0$, above which all possible cosmic trajectories with nonzero $\Omega_{\rm{m}}$ are located. The red and the green dots show the fixed points of the system that can potentially provide trajectories with desired properties, and correspond, respectively, to unstable and attractor solutions. The figure also shows a trajectory for $\lambda=1.6$ (black curve) as an indicative example—the attractor point associated with the trajectory behaves like a node. For a more complete discussion of the fixed points and their stability see Refs. Gosenca and Coles (2016); Marconnet and Tsimpis (2023); Andriot et al. (2023).
Figure 2: Phase space of the dynamical system described by eqs. 18, 19 and 20 for a hyperbolic (or open) universe (i.e., for $k=-1$). The paraboloid corresponds to the Friedmann equation (14) with $\Omega_{\rm{m}}=0$, above which all possible cosmic trajectories with nonzero $\Omega_{\rm{m}}$ are located. Trajectories that pass through the green areas experience accelerating expansion with $w_{\rm{eff}}\in[-1,-0.7]$, while the red regions correspond to $w_{\rm{eff}}=0$ (i.e., matter domination); see eq. 16. The red and green dots show the relevant fixed points of the system and correspond, respectively, to unstable and attractor solutions. For the example of $\lambda=1.6$ the attractor point behaves like a node, as shown by the indicative trajectory (black curve), where the black dot shows the initial conditions.
@@
The key point in the curvature-assisted acceleration of Refs. Marconnet and Tsimpis (2023); Andriot et al. (2023) is that the number of $e$-folds can also diverge when $1+3w_{\rm{eff}}$ approaches zero. As noted already in that reference, this is the boundary of the accelerating region—the trajectories found in Refs. Marconnet and Tsimpis (2023); Andriot et al. (2023) with curvature asymptote to fixed points living in this boundary. Although the points themselves do not support accelerated expansion, a trajectory approaching it can support an arbitrarily large number of $e$-folds. For instance, for a fixed point at $z=z_{*}$, assuming $1+3w_{\rm{eff}}\sim C(z-z_{*})$ in the vicinity of the fixed point, where $C$ is some constant, leads to $N\sim C^{-1}\log(|z-z_{*}|)$.
@@
The curvature-assisted acceleration with steep potentials is also in serious trouble when describing an inflationary phase in the early Universe. For single-field, slow-roll cosmic inflation, there is an approximate relation Baumann (2018),
@@
Furthermore, the fixed points with $\lambda>\sqrt{2}$ that lead to an arbitrary large number of $e$-folds of accelerated expansion with curvature in Refs. Marconnet and Tsimpis (2023); Andriot et al. (2023) have $z_{*}=-\Omega_{k}$ of $\mathcal{O}(1)$ unless $\lambda$ is very close to $\sqrt{2}$. That means that for large values of $\lambda$, curvature is an $\mathcal{O}(1)$ part of the energy budget of the Universe during the eternally accelerated trajectories. This contrasts with observations, either in inflationary or current era, which pose stringent upper bounds on $\Omega_{k}$.
@@ 4.1 Numerical grid analysis of phase space
Here, we solve the system of differential equations (18)-(20) for different initial conditions, i.e., for different initial values of the three phase space variables $\{x,y,z\}$, and then study the behavior of the cosmological background solution in each case. We do this through a fine grid scan of the entire phase space contained within the paraboloid of fig. 2, defined as $x^{2}+y^{2}-z\leq 1$; see eq. 14. We compute all the trajectories for the different sets of initial conditions and keep only the ones that provide an epoch of $w_{\rm{eff}}\leq-0.7$, i.e., solutions which go through the green regions of fig. 2—we show one example of such solutions for $\lambda=1.6$ in the figure (black curve). Fig. 3 presents the time evolution of $w_{\rm{eff}}$ for all these accelerating solutions. It is clear from the figure that none of the solutions exhibit an epoch of matter domination in the past, i.e., before the accelerating phase of cosmic history. In other words, none of the solutions stay within the red regions of fig. 2 for a sufficiently long time.
@@
Having said that, we do find solutions that provide extended matter domination periods, as we have demonstrated in the upper panel of fig. 4 for $\lambda=1.6$ as an example—see how $\Omega_{\rm{m}}$ behaves with time (or number of $e$-folds). However, the problem with these solutions is that none of them provide an effective equation of state parameter $w_{\rm{eff}}$ less than $\sim-0.5$ at the present time (i.e., when $\Omega_{\rm{m}}\sim 0.3$). This means that we do not obtain sufficient acceleration in these cases where matter domination exists. In the lower panel of fig. 4, we present three examples of $w_{\rm{eff}}$ evolution for the curvature-assisted quintessence model, i.e., for $\lambda=\{0.5,1.4,1.6\}$, and compare them with that of a flat $\Lambda$CDM model (i.e., with $\Omega_{k}$ set to zero) that provides a good fit to the data. The figure shows that by increasing $\lambda$ it becomes more and more difficult for the quintessence model to provide a cosmic history that is compatible with cosmological observations, as the current value of $w_{\rm{eff}}$ does not reach the observationally preferred value of $\sim-0.7$, as indicated by our $\Lambda$CDM example.
@@ 4.2 Bayesian analysis and parameter estimation
We perform a Markov chain Monte Carlo (MCMC) scan of the parameter space of the quintessence model for two cases. [^3] The Mathematica codes used in our analysis will be made publicly available in a GitHub repository upon publication of the paper. We first consider a flat universe, i.e., we fix the present value of the curvature density parameter $\Omega_{k,0}$ to zero. We then allow $\Omega_{k,0}$ to vary as a free parameter. For each case, we perform two separate analyses, one by setting the initial value of the scalar field’s time derivative $\dot{\phi}_{0}$ to zero and one by varying it as a free parameter. In both cases, we fix $\phi_{0}$ (the initial value of the scalar field) to zero and vary the parameter $V_{0}$ in the potential $V(\phi)=V_{0}e^{-\lambda\phi}$, as one of the two parameters is redundant and can be replaced by the other. We impose flat (or uniform) priors on all free parameters—table 1 shows the ranges of parameters we employ in our scans. Comparing the parameter constraints we obtain for different cases will tell us whether assisting the accelerated expansion with a nonzero spatial curvature will increase the upper bound on the parameter $\lambda$ so that large values suggested by the statement (2) become observationally viable.
@@
$\Omega_{\rm{m,0}}$ $[0.001,1]$
$\omega_{\mathrm{b,0}}$ $[0.01,0.04]$
$\Omega_{k,0}$ $[-0.2,0.2]$
@@
Table 1: Ranges of model parameters employed in our scans. Note that $\dot{\phi}_{0}$ is fixed to zero in some of the analyses, so the presented range for $\dot{\phi}_{0}$ is only for cases where it is varied as a free parameter. $\Omega_{\rm{m,0}}$, $\omega_{\mathrm{b,0}}$, and $\Omega_{k,0}$ are the present values of $\Omega_{\rm{m}}$, $\omega_{\mathrm{b}}$, and $\Omega_{k}$, respectively.
Figure 5: The $68.3\%$ and $95.5\%$ confidence regions and one-dimensional marginalized probability density functions for the quintessence model of dark energy with an exponential potential when the spatial curvature of the Universe is set to zero (i.e., $\Omega_{k,0}=0$, where $\Omega_{k,0}$ is the present value of the curvature density parameter $\Omega_{k}$). We show the results for two cases: 1) the initial value of the scalar field’s time derivative $\dot{\phi}_{0}$ is set to zero, and 2) $\dot{\phi}_{0}$ is varied as a free parameter.
Figure 6: As in fig. 5, but for cases where the spatial curvature of the Universe is allowed to be nonzero, i.e., the present value of the curvature density parameter, $\Omega_{k,0}$, is varied as a free parameter.
 Parameter   $\Omega_{k,0}=0$ (flat universe), $\dot{\phi}_{0}=0$   $\Omega_{k,0}=0$ (flat universe), free $\dot{\phi}_{0}$   Free $\Omega_{k,0}$, $\dot{\phi}_{0}=0$   Free $\Omega_{k,0}$, free $\dot{\phi}_{0}$
$\Omega_{\rm{m,0}}$ $0.3050^{+0.0784}_{-0.0591}$ $0.3269^{+0.0730}_{-0.0777}$ $0.3134^{+0.0849}_{-0.0441}$ $0.3699^{+0.0820}_{-0.0900}$
$\omega_{\mathrm{b,0}}$ $0.0226^{+0.0002}_{-0.0001}$ $0.0226^{+0.0002}_{-0.0001}$ $0.0225^{+0.0001}_{-0.0002}$ $0.0224^{+0.0002}_{-0.0001}$
$\Omega_{k,0}$ N/A N/A $0.0047^{+0.0018}_{-0.0023}$ $0.0056^{+0.002}_{-0.003}$
@@
We present the results of our MCMC analysis in figs. 5 and 6, where two-dimensional $68.3\%$ and $95.5\%$ confidence regions and one-dimensional marginalized probability density functions (PDFs) for different parameters and different cases are provided. We also present in table 2 $1\sigma$ (i.e., $68.3\%$ confidence level) constraints on the parameters. Since the parameter of interest here is $\lambda$, let us compare our constraints on that parameter as given in the figures and the table for the different cases we have studied. Fig. 5 and table 2 show that cosmological observations place an upper bound of $\sim 0.6$ on $\lambda$ when $\dot{\phi}_{0}$ is fixed to zero, and an upper bound of $\sim 0.5$ when $\dot{\phi}_{0}$ is varied, at the $63.3\%$ confidence level, which are consistent with the results of Refs. Agrawal et al. (2018); Akrami et al. (2019); Raveri et al. (2019). When $\Omega_{k,0}$ is allowed to vary, we obtain the upper bound of $\sim 0.7$ on $\lambda$ (as given in fig. 6 and table 2), independently of whether $\dot{\phi}_{0}$ is varied or fixed to zero, even though the contours and the one-dimensional PDFs in the two cases are slightly different. Note, however, that it is important to vary $\dot{\phi}_{0}$ in the case of free $\Omega_{k,0}$ for full exploration of the phase space of the quintessence model—there is no a priori reason for setting it to zero. It is, therefore, true that allowing $\Omega_{k,0}$ to vary pushes the upper bound on $\lambda$ to larger values, but the improvement is tiny and the favored values of $\lambda$ are far below the $\lambda=\sqrt{2}$ limit of the statement (2). This means that the new fixed point in the phase space of the model which arises as a result of $\Omega_{k}$ taking positive values, i.e., the Universe having a negative curvature, does not help with making $\lambda>\sqrt{2}$ consistent with observations—this is what we expected from the qualitative analyses of previous sections. It is, however, interesting to note that when $\Omega_{k,0}$ is allowed to vary, cosmological data seem to prefer nonzero curvature and therefore indicate a slight preference for an open universe, although the preference is only slightly more than $2\sigma$. We do not perform a model comparison here, but it is likely that the slightly larger upper bound on $\lambda$ for the curvature-assisted model compared to the flat one is due to the increase in the number of free parameters (i.e., the number of degrees of freedom) in the former case. As stated above, the results of the full MCMC analysis perfectly agree with the findings of the numerical grid analysis of section 4.1.